# Constructing a T-test for Value Function Comparison of Individualized Treatment Regimes in the Presence of Multiple Imputation for Missing Data

### v1

@@ Appendix A Proofs
in distribution, and standard arguments can now be used to show that $\hat{T}_{m}^{2}\rightarrow T_{0}^{2}$ as $m\rightarrow\infty$, where
$\hat{T}_{m}^{2}=m^{-1}\sum_{i=1}^{m}(\hat{U}_{i,1}-\hat{U}_{i,2}-\bar{U}_{m,1}+\bar{U}_{m,2})^{2}+m(\hat{W}_{m,1}-\hat{W}_{m,2})^{T}\hat{\Sigma}_{n}(\hat{W}_{m,1}-\hat{W}_{m,2}).$
Appendix B Additional Figures
@@
Figure 5: The value function results for both training and test sets for five different treatment regimes: observed treatment, treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, Q-learning optimal ITR, and D-learning optimal ITR. For comparison purposes, we include the average outcome as a benchmark. The results on original data (a) and three simulated data sets with three levels of treatment effect modification are shown (b) $\delta=1$, (c) $\delta=2$, (d) $\delta=10$.
@@
Figure 6: The differences of value functions between the pairs of the following treatments for both training and test sets: treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, Q-learning optimal ITR, and D-learning optimal ITR. The results on original data (a) and three simulated data with three levels of treatment effect modification are shown (b) $\delta=1$, (c) $\delta=2$, (d) $\delta=10$.

### v3

@@ Appendix A Proofs
and standard arguments can now be used to show that $\hat{T}_{m}^{2}\rightarrow T_{0}^{2}$ as $m\rightarrow\infty$, where
$\hat{T}_{m}^{2}=m^{-1}\sum_{i=1}^{m}(\hat{U}_{i,1}-\hat{U}_{i,2}-\bar{U}_{1}+\bar{U}_{2})^{2}+\frac{m}{n}(\hat{W}_{1}-\hat{W}_{2})^{T}\hat{\Sigma}_{n}(\hat{W}_{1}-\hat{W}_{2}).$
Proof of Assumptions: logistic regression as an example
Proof.
In general, most generalized linear models satisfy all assumptions under standard regularity conditions. Here, we illustrate this using logistic regression as an example. The first assumption is automatically satisfied for logistic regression. Suppose the propensity score model is specified as $\pi(a|x,\theta_{0})=\frac{e^{a\theta_{0}^{T}x}}{1+e^{\theta_{0}^{T}x}}$, for $a\in\{0,1\}$.
The left-hand side of Equation (1) becomes
$\sqrt{n}\left(\frac{e^{a\hat{\theta}_{n}^{T}x}}{1+e^{\hat{\theta}_{n}^{T}x}}-\frac{e^{a\theta_{0}^{T}x}}{1+e^{\theta_{0}^{T}x}}\right).$
Using a Taylor expansion of $\hat{\theta}_{n}$ around $\theta_{0}$ for both $a=1$ and $a=0$, we obtain:
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})^{T}(2a-1)x\cdot\frac{e^{a\theta_{0}^{T}x}}{(1+e^{\theta_{0}^{T}x})^{2}}+o_{P}(1).$
Therefore, we define
$\phi_{0}(a,x)=x(2a-1)\cdot\frac{e^{a\theta_{0}^{T}x}}{(1+e^{\theta_{0}^{T}x})^{2}},\quad\text{and}\quad\hat{\phi}_{n}(a,x)=x(2a-1)\cdot\frac{e^{a\hat{\theta}_{n}^{T}x}}{(1+e^{\hat{\theta}_{n}^{T}x})^{2}}.$
This confirms that Assumption 1 is satisfied.
Assumption 2 holds by the asymptotic normality of the maximum likelihood estimator (MLE) for logistic regression, provided the standard covariance estimator from the MLE is used. An implication of this assumption is that the propensity score model must be correctly specified.
We can also verify that the third assumption holds. From Assumption 2, we have that $\hat{\theta}_{n}\rightarrow\theta_{0}$ in probability as $n\rightarrow\infty$. By the continuous mapping theorem, it follows that $\hat{\phi}_{n}(A,X)\rightarrow\phi_{0}(A,X)$ in probability as $n\rightarrow\infty$.
Moreover, since $\|\hat{\phi}_{n}(A,X)-\phi_{0}(A,X)\|\leq\|X\|$ and $E\|X\|^{2}<\infty$, we can apply the dominated convergence theorem to conclude that
$P\|\hat{\phi}_{n}(A,X)-\phi_{0}(A,X)\|^{2}\rightarrow 0,$
in probability as $n\rightarrow\infty$.
Appendix B Additional Tables and Figures
Table S1: Estimated value functions of different treatment regimes for CHNS data.
Training Set Test Set
Mean SD Mean SD
Obs 0.602 0.021 $0.595$ 0.031
AllLowPA 0.603 0.018 $0.591$ 0.026
AllHighPA 0.600 0.011 $0.598$ 0.017
Q-Learning 0.574 0.013 $0.573$ 0.017
D-Learning 0.576 0.016 $0.598$ 0.023
Table S2: Estimated value differences and p-values of selected comparisons between treatment regimes for CHNS data.
Training Set Test Set
Difference Mean SD P-Value Mean SD P-Value
AllLowPA vs. Obs 0.001 0.011 0.924 -0.004 0.017 0.827
AllHighPA vs. Obs -0.001 0.018 0.953 0.003 0.026 0.897
AllHighPA vs. AllLowPA -0.002 0.021 0.921 0.007 0.032 0.821
Q vs. Obs -0.028 0.017 0.096 -0.022 0.026 0.407
Q vs. AllLowPA -0.029 0.017 0.086 -0.018 0.026 0.497
Q vs. AllHighPA -0.027 0.013 0.043 -0.025 0.018 0.156
Q vs. D -0.002 0.013 0.871 -0.024 0.020 0.223
Figure S1: Value functions and 95% confidence interval estimates obtained using propensity scores obtained from different propensity score models: (a) a logistic regression model including all covariates $\hat{\pi}=\hat{\theta}X$, (b) a correctly specified model containing only truly relevant covariates with estimated parameters $\hat{\pi}=\hat{\theta}^{\top}[1,X_{1},X_{2}]$, (c) the true propensity score model with known parameters $\pi_{0}=\theta_{0}^{\top}[1,X_{1},X_{2}]$, and (d) an incorrectly specified model including only non-relevant covariates with estimated parameters $\hat{\pi}=\hat{\theta}^{\top}[1,X_{3},X_{10},X_{15},X_{20}]$. Results shown here are for scenario 4 (observational data with data missing at random) with $n=5000$ training samples, and $\delta_{t}=1$. Scenario 3 shows similar results and is thus omitted. When the propensity score is incorrect, the true value function remains the same for $d_{1}$ $d_{2}$ $d_{opt}$, the estimates have larger bias, and thus the true variance becomes larger. The true variance for value difference equals the average of the squared difference between the estimated value difference and the true value difference.
Figure S2: Differences between the two value functions and their corresponding 95% confidence interval estimates, obtained using propensity scores derived from different models: (a) a logistic regression model including all covariates $\hat{\pi}=\hat{\theta}X$, (b) a correctly specified model containing only truly relevant covariates with estimated parameters $\hat{\pi}=\hat{\theta}^{\top}[1,X_{1},X_{2}]$, (c) the true propensity score model with known parameters $\pi_{0}=\theta_{0}^{\top}[1,X_{1},X_{2}]$, and (d) an incorrectly specified model including only non-relevant covariates with estimated parameters $\hat{\pi}=\hat{\theta}^{\top}[1,X_{3},X_{10},X_{15},X_{20}]$. Results shown here are for scenario 4 with n=5000 training samples and $\delta_{t}=1$. Scenario 3 shows similar results and is thus omitted.
Figure S3: A demonstration of one estimated propensity score compared to the true propensity score in propensity score model misspecification sensitivity analysis in one replication.
@@ Appendix B Additional Tables and Figures
Figure S4: The value function results for both training and test sets for five different treatment regimes: observed treatment, treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, Q-learning optimal ITR, and D-learning optimal ITR. For comparison purposes, we include the average outcome as a benchmark. The results on original data (a) and three simulated data sets with three levels of treatment effect modification are shown (b) $\delta=1$, (c) $\delta=2$, (d) $\delta=10$.
@@
Figure S5: The differences of value functions between the pairs of the following treatments for both training and test sets: treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, Q-learning optimal ITR, and D-learning optimal ITR. The results on original data (a) and three simulated data with three levels of treatment effect modification are shown (b) $\delta=1$, (c) $\delta=2$, (d) $\delta=10$.